Melanie's weight is equal to the weight of 20 rabbits.The weight of 50 rabbits is equal to the weight of 3 goats.The weight of 15 goats is equal to the weight of 12 horses.The weight of 20 horses is equal to the weight of 12 cows.
The weight of 60 goats is equal to how many times the weight of Melanie.?????????????The weight of 24 horeses is equal to how many times the weight of Melanie???????The weight of 48 cows is equal to how many times the weight of melanie???PLZZZZZZZZZZ ANSWERRRRRRRRRRRRR HELPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP .FOR GODDDDDDDDD SAKEEEEEEEEE !!!!!!!!!!!!!!!!!!! VERY IMPORTANT !!!!!!!!!!

Answer by eperette(173)

GIVEN
M = 20R
50R = 3G
15G = 12H
20H = 12C
*************
Unknown:
60G = ? M
24H = ? M
48C = ? M
****************
60G * (50R/3G) * (1M/20R) = 50M
24H * (15G/12H) * (50R/ 3G)* (1M/20R) = 25M
48C * (20H/12C) * (15G/12H)* (50R/3G) * (1M/20R) = 83.3333M